### v1

@@ Abstract
Soil quality (SQ) plays a crucial role in sustainable agriculture, environmental conservation, and land-use planning. Traditional SQ assessment techniques rely on costly, labor-intensive sampling and laboratory analysis, limiting their spatial and temporal coverage. Advances in Geographic Information Systems (GIS), remote sensing, and machine learning (ML) enabled efficient SQ evaluation. This paper presents a comprehensive roadmap distinguishing it from previous reviews by proposing a unified and modular pipeline that integrates multi-source soil data, GIS and remote sensing tools, and machine learning techniques to support transparent and scalable soil quality assessment. It also includes practical applications. Contrary to existing studies that predominantly target isolated soil parameters or specific modeling methodologies, this approach consolidates recent advancements in Geographic Information Systems (GIS), remote sensing technologies, and machine learning algorithms within the entire soil quality assessment pipeline. It also addresses existing challenges and limitations while exploring future developments and emerging trends in the field that can deliver the next generation of soil quality systems making them more transparent, adaptive, and aligned with sustainable land management.
@@
Soil Quality Analysis, GeoSpatial Analysis, Geographic Information Systems, roadmap.
@@ 1 Introduction
Soil quality plays a fundamental role in agriculture, environmental sustainability, and land management. It supports food security, ecosystem resilience, and human well-being [sumathi_improved_2023]. The concept of soil quality was initially formalized in the early 1990s, refers to the capacity of soil to perform essential tasks. According to the USDA (1994) [^1] USDA: United States Department of Agriculture, Soil quality is defined as the capacity of a given soil type, within its inherent or managed ecological context, to preserve or enhance air and water quality, sustain biotic productivity, and support the health of both human populations and ecological systems [Swapana_Sepehya2024].
@@
The remainder of this paper is structured as follows : in Section 2 we describe soil properties and indices. Section 3 is dedicated to the basics of geospatial analysis systems for soil quality analysis, followed by the tools and frameworks available in 4. Section 5 reviews related work. Section 6 discusses key challenges and limitations. Followed by the trends and future research directions highlighted in Section 7. Finally, Section 8 presents the conclusion.
@@ 2 Soil quality characteristics and indicators
The key physical indicators are soil texture, bulk density, porosity, aggregate stability, and infiltration rate. Soil texture affects water retention, nutrient holding capacity, and root penetration. While bulk density is an indicator of compaction and aeration status. These parameters are fundamental in assessing soil structure and hydrological behavior, and can be modeled using Digital Soil Mapping (DSM) techniques combined with terrain attributes derived from Digital Elevation Models (DEMs).
@@ 3 Geospatial systems for soil quality analysis : proposed Pipeline
A fundamental concept in spatial analysis is the Geographic Information System GIS, which is defined as "a computer-based system for the acquisition, storage, management, querying, analysis, and visualization of geospatial data" [changGISintroduction2019]. It was introduced by Dr. Roger Tomlinson, known as the father of GIS, who developed the first GIS system known as the Canada Geographic Information System (CGIS), which was used to catalog and analyze Canada’s natural resources [lawheadPyHist2019].
Based on the reviewed literature, this study introduces a unified, modular GIS-based soil quality analysis pipeline, as illustrated in Fig. 1. Guided by existing research [changGISintroduction2019, lawheadPyHist2019, lovelaceGeocomputationR2020], the proposed integrated architecture and workflow constitute an innovative consolidation. The pipeline synthesizes heterogeneous data sources, computational tools, and analytical methodologies into a cohesive system designed for scalable, transparent, and adaptive soil quality evaluation. It functions as a reference architecture to facilitate the development and enhancement of geospatial soil monitoring systems by researchers and practitioners.
Figure 1: GIS analysis pipeline system.
@@ 3.1 Data collection and Storage
Figure. 2 illustrates the main spatial object types namely points, lines, and polygons [lovelaceGeocomputationR2020], that constitute the fundamental geometric components of geospatial data models employed in soil analysis.
@@
Precise comprehension of these geometric primitives is critical for the organization of soil datasets and the implementation of suitable spatial analysis procedures within Geographic Information Systems (GIS). These objects can be stored as Vectors or Raster, as illustrated by Figure 3, where :
  • Vectors that contains the information about points, lines or polygons, with it specific coordinates.
@@
  • AfSIS [^2] https://www.isric.org/projects/africa-soil-information-service-afsis Soil datasets and maps for Africa. They focus specifically on sub-saharan Africa, which maps soil properties such as: soil organic carbon, pH, texture (sand, silt, clay), exchangeable nutrients (Ca, Mg, K, etc.), cation exchange capacity (CEC), bulk density.
@@
  • FAO Global Soil Map [^5] https://www.fao.org/soils-portal/data-hub/soil-maps-and-databases/en/ is a collection of global and regional soil information products developed and maintained by the Food and Agriculture Organization (FAO). It contains GSOCmaps (Global Soil Organic Carbon Map), GloSIS (Global Soil Information System), HWSD (Harmonized World Soil Database), soil pH, texture, salinity, nutrient status properties.
@@
  • Soil Data Mart (USDA NRCS) [^7] https://websoilsurvey.nrcs.usda.gov/ developed by the United States Department of Agriculture’s Natural Resources Conservation Service (USDA NRCS) to provide access to official soil survey data across the United States. It has now been replaced by the Web Soil Survey (WSS) and integrated into the Soil Data Access (SDA) system, which continues to provide public access to NRCS-certified soil data and contains soil data: texture, structure, and horizon data, soil organic matter and pH, Hydrologic group and drainage, land capability classes, engineering and building site development interpretations, crop suitability and forest productivity.
  • MODIS Soil Moisture Data (NASA) [^8] https://modis.gsfc.nasa.gov/data/ based on data from MODIS instruments aboard NASA’s Terra (since 1999) and Aqua (since 2002) satellites. It provide information about soil moisture which is not directly sensed but modeled using indices derived from MODIS, such as: NDVI/EVI (vegetation health), LST (land surface temperature), Albedo, emissivity.
@@
  • ISRIC - World Soil Information [^10] https://www.isric.org/ is a global soil data platform developed and maintained by International Soil Reference and Information Centre (SRIC). It provides access to a harmonized, curated, and geo-referenced database of soil profile observations from around the world. It offers consistent soil point data for various soil properties : soil organic carbon (SOC), pH (H2O and KCl), texture (sand, silt, clay), bulk density, cation exchange capacity (CEC), base saturation, Electrical conductivity (EC), Nutrients (e.g., P, K, N).
@@
  • LandPKS Soil Data [^12] https://landpks.org/ combines field observations with soil science knowledge to assess land potential, soil health, and guide sustainable land management decisions. It empowers users with tools to gather soil profile data directly in the field using smartphones, helping fill data gaps in underserved regions worldwide. Th field-measured or observed: Soil texture by feel (sand, silt, clay), Soil color (using Munsell charts), Soil depth and rooting depth, Rock fragment content, Surface and subsurface horizon characteristics.
@@
The comparison presented in Table 1 delineates the trade-offs among spatial extent, spatial resolution, and data typology across principal soil datasets. Globally distributed, model-generated datasets such as SoilGrids and MODIS offer lower spatial resolution, making them more suitable for macro-scale spatial analyses. Conversely, in-situ datasets like LUCAS and LandPKS deliver higher measurement precision at the field scale, making them well-suited for localized studies and validation of predictive models. Layered or multi-parameter systems such as ESDB and SIS are optimized for integration within GIS frameworks and thematic mapping applications. Specialized datasets, including SWAT, are tailored for hydrological modeling purposes. Selection of an appropriate dataset must consider the spatial analytical scale, specific research requirements, and the intended modeling approach.
@@ 3.2 Preprocessing
Preprocessing is a essential stage in geospatial soil quality analysis systems, acting as a crucial intermediary between data collection and effective analytical modeling stages. Given the variety of data formats, spatial resolutions, and coordinate systems, preprocessing ensures that datasets are accurately cleaned, standardized, and spatially aligned. This process facilitates the integration of multiple sources of geospatial information. The Geographic Information System (GIS) operations include [lawheadPyHist2019] :
@@
Since the Earth has a spherical shape, projection is a way to represent the coordinates of an object using a coordinate reference system (CRS). This representation preserves common properties, such as area, scale, bearing, distance, or shape. Projections are commonly delineated by a collection of more than 40 parameters, presented in either XML or a textual form known as Well-Known Text (WKT), which serves to specify the transformation algorithm. Those WKT are registered by the International Association of Oil and Gas Producers (IOGP) and were formerly known as the European Petroleum Survey Group (EPSG), representing over 5000 entries. For example, the Google Mercator projection used in Google Maps utilizes the code EPSG:3857.
@@
Figure 4: GIS overly operations
@@ 3.3 Analysis
The aim of Machine Learning (ML) is to establish computational models that can automatically learn patterns and generalize from input data [PeterflachML2012]. It offers a collection of algorithms that can handle various types of data (labeled, unlabeled; unstructured, structured, semi-structured; sequential, unsequential) to address crucial challenges: classification, regression, clustering, dimensionality reducing, sequential analysis.
@@
For spatial interpolation, an alternative neural network architecture designated as FBRN is utilized, wherein the activation functions of the hidden layers are Radial Basis Functions (RBFs), originally introduced by Powell (1987) to address multivariate interpolation problems [linChenRBFN2004]. Furthermore, in 2017, a Convolutional Neural Network (CNN) architecture was utilized by A.U. Waldeland and A. Solberg for image-based seismic interpretation [70YearMLinGeo2020].
@@ 3.4 Evaluation metrics
  • R2 (Coefficient of Determination) : Indicates how well the model explains the variability in the observed data. A closer to $1$ values of R2 means a better fit. It is given by the formula 2.
@@
  • Accuracy It’s the most basic performance metric, which represent the proportion of correct predictions over total predictions. It is computed using the formula 4.
@@
  • Precision represents the proportion of true positive predictions out of all positive predictions made. It is used when false positives are costly, it’s formula is given by 5.
@@
Where TP is True Positive, the number of items that are classified in its proper class; TN is True Negative represent cases where the model accurately predicts the absence of a class; FP is False Positive, it occur when the model incorrectly classifies a negative case as positive, leading to a Type I error. and FN extends for False Negative appers when the model fails to detect a condition that is actually present, constituting a Type II error.
@@ 4 Available tools and frameworks
  • Sentinel satellites, part of the Copernicus program, provides high-resolution imagery (10m spatial resolution) across 13 spectral bands, from visible to shortwave infrared.
  • MODIS (Moderate Resolution Imaging Spectroradiometer), aboard NASA’s Terra and Aqua satellites, offers daily global coverage and provides data at a moderate spatial resolution (250m to 1km).
@@
The creation of general and multipurpose systems required a complex and large scale data from different sources and computational resources. Geospatial data analysis can be performed using Big data tools, such as Hadoop-based frameworks, such as SpatialHadoop [^17] https://spatialhadoop.cs.umn.edu/ and HadoopGIS [^18] https://github.com/bunnyg/Hadoop-GIS, or Spark-based frameworks, including GeoTrellis [^19] http://geotrellis.github.io/ and GeoSpark [^20] https://github.com/agile-lab-dev/GeoSpark. Additionally, other proposed solutions, such as BiGeo[Liu2019], may also be utilized depending on the specific application requirements.
@@ 5 Applications
  • Climate Interaction & Adaptation : explores how soil data supports climate modeling and mitigation.
@@ 5.1 Precision Agriculture
Ramzan et al. [Ramzan2024] proposed a two-module crop recommendation system, integrating environmental and soil parameters to guide crop selection. The first module utilizes a combination of static data, such as publicly available datasets and machine learning algorithms including K-Nearest Neighbors (KNN), Decision Tree, Random Forest, Support Vector Machine (SVM), LightGBM, CatBoost, and AdaBoost. This module was trained on the Kaggle “Crop Recommendation Dataset”, which contains approximately 100 samples for each of 22 crop types. The second module incorporates real-time data from IoT sensors deployed in agricultural areas of Punjab, Pakistan, focusing on six crops (wheat, rice, cotton, maize, gram, and groundnut), supplemented by manual inputs from local farmers.
@@ 5.2 Soil Fertility Analysis
Furthermore, [folorunso_exploring_2023] examined supervised machine learning algorithms, specifically, Random Forest (RF), Support Vector Machines (SVM), and Deep Neural Networks (DNN) for quantitative prediction of soil nutrient properties within precision agriculture. The research emphasized advancements in Digital Soil Mapping (DSM) methodologies and demonstrated the integration of Diffuse Reflectance Infrared Spectroscopy (DRIS) with Convolutional Neural Networks (CNN) for spectral data analysis, applying this combined approach to the LUCAS soil dataset to quantitatively estimate six soil chemical parameters.
On the other hand, [el_behairy_accurate_2024] developed a Soil Quality Index (SQI) model utilizing 16 physical, chemical, and fertility soil attributes derived from laboratory analyses, employing an Artificial Neural Network (ANN) implemented in MATLAB, without incorporating geospatial data.
In study [mine_comparative_2025], authors compared four machine learning algorithms to identify the most effective regressor to predict soil quality index using the SoilGrid dataset across four European countries. The selected model was subsequently utilized to generate soil quality maps.

### v2

@@ Abstract
Soil quality plays a crucial role in sustainable agriculture, environmental conservation, and land-use planning. Traditional SQ assessment techniques rely on costly, labor-intensive sampling and laboratory analysis, limiting their spatial and temporal coverage. Advances in Geographic Information Systems (GIS), remote sensing, and machine learning (ML) enabled efficient soil quality evaluation. This paper presents a comprehensive roadmap distinguishing it from previous reviews by proposing a unified and modular pipeline that integrates multi-source soil data, GIS and remote sensing tools, and machine learning techniques to support transparent and scalable soil quality assessment. It also includes practical applications. Contrary to existing studies that predominantly target isolated soil parameters or specific modeling methodologies, this approach consolidates recent advancements in Geographic Information Systems (GIS), remote sensing technologies, and machine learning algorithms within the entire soil quality assessment pipeline. It also addresses existing challenges and limitations while exploring future developments and emerging trends in the field that can deliver the next generation of soil quality systems making them more transparent, adaptive, and aligned with sustainable land management.
@@
Soil Quality Assessment, GeoSpatial Analysis, Geographic Information Systems, Roadmap, Machine Learning, Sustainable Agriculture; Land Management.
@@ 1 Introduction
Soil quality plays a fundamental role in agriculture, environmental sustainability, and land management. It supports food security, ecosystem resilience, and human well-being [sumathi_improved_2023]. The concept of soil quality was initially formalized in the early 1990s, refers to the capacity of soil to perform essential tasks. According to the definition given by USDA (United States Department of Agriculture) in 1994, soil quality is defined as the capacity of a given soil type, within its inherent or managed ecological context, to preserve or enhance air and water quality, sustain biotic productivity, and support the health of both human populations and ecological systems [Swapana_Sepehya2024].
@@
The remainder of this paper is structured as follows : in Section 2 we presented a foundational understanding of soil quality properties and indices. Section 3 is dedicated to the basics of geospatial analysis systems for soil quality analysis, followed by the tools and frameworks available in 4. Section 5 reviews related work. Section 6 discusses key challenges and limitations. Followed by the trends and future research directions highlighted in Section 7. Finally, Section 8 presents the conclusion.
@@ 2 Soil quality characteristics and indicators
The key physical indicators are soil texture, bulk density, porosity, aggregate stability, and infiltration rate. The texture of the soil affects the retention of water, the holding capacity of nutrients and the penetration of the roots. While bulk density is an indicator of compaction and aeration status. These parameters are fundamental in assessing soil structure and hydrological behavior, and can be modeled using Digital Soil Mapping (DSM) techniques combined with terrain attributes derived from Digital Elevation Models (DEMs).
@@ 3 Geospatial Systems for Soil Quality Analysis: Proposed Pipeline
A fundamental concept in spatial analysis is GIS, which is defined as "a computer-based system for the acquisition, storage, management, querying, analysis, and visualization of geospatial data" [changGISintroduction2019]. It was introduced by Dr. Roger Tomlinson, known as the father of GIS, who developed the first GIS system known as the Canada Geographic Information System (CGIS), which was used to catalog and analyze Canada’s natural resources [lawheadPyHist2019].
Based on the reviewed literature, this study introduces a unified, modular GIS-based soil quality analysis pipeline, as illustrated in Fig. 1. Based on existing research [changGISintroduction2019, lawheadPyHist2019, lovelaceGeocomputationR2020], the proposed integrated architecture and workflow constitute an innovative consolidation. The pipeline synthesizes heterogeneous data sources, computational tools, and analytical methodologies into a cohesive system designed for scalable, transparent, and adaptive soil quality evaluation. It functions as a reference architecture to facilitate the development and enhancement of geospatial soil monitoring systems by researchers and practitioners.
Figure 1: Proposed GIS analysis pipeline system.
@@ 3.1 Data collection and Storage
Figure. 2 illustrates the main types of GIS spatial objects, namely points, lines, and polygons [lovelaceGeocomputationR2020], that constitute the fundamental geometric components of the geospatial data models employed in soil analysis.
@@
Precise comprehension of these geometric primitives is critical for the organization of soil datasets and the implementation of suitable spatial analysis procedures within GIS. These objects can be stored as Vectors or Raster, as illustrated by Figure 3, where :
  • Vectors that contain the information about points, lines or polygons, with their specific coordinates.
@@
  • AfSIS [^1] https://www.isric.org/projects/africa-soil-information-service-afsis Soil datasets and maps for Africa. It focuses specifically on sub-saharan Africa, which maps soil properties such as: soil organic carbon, pH, texture (sand, silt, clay), exchangeable nutrients (Ca, Mg, K, etc.), CEC, bulk density.
@@
  • FAO (Food and Agriculture Organization) Global Soil Map [^4] https://www.fao.org/soils-portal/data-hub/soil-maps-and-databases/en/ is a collection of global and regional soil information products developed and maintained by the FAO Organization. It contains GSOCmaps (Global Soil Organic Carbon Map), GloSIS (Global Soil Information System), HWSD (Harmonized World Soil Database), soil pH, texture, salinity, nutrient status properties.
@@
  • Soil Data Mart (USDA NRCS) [^6] https://websoilsurvey.nrcs.usda.gov/ developed by the USDA’s Natural Resources Conservation Service (USDA NRCS) to provide access to official soil survey data across the United States. It has now been replaced by the Web Soil Survey (WSS) and integrated into the Soil Data Access (SDA) system, which continues to provide public access to NRCS-certified soil data and contains soil data: texture, structure, and horizon data, soil organic matter and pH, Hydrologic group and drainage, land capability classes, engineering and building site development interpretations, crop suitability and forest productivity.
  • MODIS Soil Moisture Data (NASA) [^7] https://modis.gsfc.nasa.gov/data/ based on data from MODIS instruments aboard NASA’s Terra (since 1999) and Aqua (since 2002) satellites. It provides information with spatial resolution (250m to 1km) about soil moisture which is not directly sensed but modeled using indices derived from MODIS, such as: NDVI/EVI (vegetation health), LST (land surface temperature), Albedo, emissivity.
@@
  • ISRIC - World Soil Information [^9] https://www.isric.org/ is a global soil data platform developed and maintained by International Soil Reference and Information Centre (SRIC). It provides access to a harmonized, curated, and geo-referenced database of soil profile observations from around the world. It offers consistent soil point data for various soil properties : OC, pH (H2O and KCl), texture (sand, silt, clay), bulk density, CEC, base saturation, EC, Nutrients (e.g., P, K, N).
@@
  • LandPKS Soil Data [^11] https://landpks.org/ combines field observations with soil science knowledge to assess land potential, soil health, and guide sustainable land management decisions. It empowers users with tools to gather soil profile data directly in the field using smartphones, helping fill data gaps in underserved regions worldwide. The field-measured or observed: Soil texture by feel (sand, silt, clay), Soil color (using Munsell charts), Soil depth and rooting depth, Rock fragment content, Surface and subsurface horizon characteristics.
@@
The comparison presented in Table 1 delineates the trade-offs among spatial extent, spatial resolution, and data typology across principal soil datasets. Globally distributed, model-generated datasets such as SoilGrids and MODIS offer lower spatial resolution, making them more suitable for macro-scale spatial analyses. Conversely, in-situ datasets like LUCAS and LandPKS deliver higher measurement precision at the field scale, making them well-suited for localized studies and validation of predictive models. Layered or multi-parameter systems such as European Soil Database (ESDB) and SIS are optimized for integration within GIS frameworks and thematic mapping applications. Specialized datasets, including SWAT, are tailored for hydrological modeling purposes. Selection of an appropriate dataset must consider the spatial analytical scale, specific research requirements, and the intended modeling approach.
@@ 3.2 Preprocessing
Preprocessing is an essential stage in geospatial soil quality analysis systems, acting as a crucial intermediary between data collection and effective analytical modeling stages. Given the variety of data formats, spatial resolutions, and coordinate systems, preprocessing ensures that datasets are accurately cleaned, standardized, and spatially aligned. This process facilitates the integration of multiple sources of geospatial information. The GIS operations include [lawheadPyHist2019] :
@@
Since the Earth has a spherical shape, projection is a way to represent the coordinates of an object using a coordinate reference system (CRS). This representation preserves common properties, such as area, scale, bearing, distance, or shape. Projections are commonly delineated by a collection of more than 40 parameters, presented in either XML or a textual form known as Well-Known Text (WKT), which serves to specify the transformation algorithm. Those WKT are registered by the International Association of Oil and Gas Producers (IOGP) and were formerly known as the European Petroleum Survey Group (EPSG), representing over 5000 entries. For example, the Google Mercator projection used in Google Maps uses the code EPSG:3857.
@@
Figure 4: GIS overlay operations
@@ 3.3 Analysis
The aim of ML is to establish computational models that can automatically learn patterns and generalize from input data [PeterflachML2012]. It offers a collection of algorithms that can handle various types of data (labeled, unlabeled; unstructured, structured, semi-structured; sequential, unsequential) to address crucial challenges: classification, regression, clustering, dimensionality reducing, sequential analysis.
@@
For spatial interpolation, an alternative neural network architecture designated as Radial Basis Functions Network (RBFN) is utilized, wherein the activation functions of the hidden layers are Radial Basis Functions (RBFs), originally introduced by Powell (1987) to address multivariate interpolation problems [linChenRBFN2004]. Furthermore, in 2017, a Convolutional Neural Network (CNN) architecture was utilized by A.U. Waldeland and A. Solberg for image-based seismic interpretation [70YearMLinGeo2020].
@@ 3.4 Evaluation metrics
  • R2 (Coefficient of Determination) : Indicates how well the model explains the variability in the observed data. A closer to $1$ values of R2 mean a better fit. It is given by the formula 2.
@@
  • Accuracy It’s the most basic performance metric, which represents the proportion of correct predictions over total predictions. It is computed using the formula 4.
@@
  • Precision represents the proportion of true positive predictions out of all positive predictions made. It is used when false positives are costly, its formula is given by 5.
@@
Where TP is True Positive, the number of items that are classified in its proper class; TN is True Negative represents cases where the model accurately predicts the absence of a class; FP is False Positive occurs when the model incorrectly classifies a negative case as positive, leading to a Type I error. and FN extends for False Negative appears when the model fails to detect a condition that is actually present, constituting a Type II error.
@@ 4 Available tools and frameworks
  • Sentinel satellites, part of the Copernicus program, provide high-resolution imagery (10m spatial resolution) across 13 spectral bands, from visible to shortwave infrared.
@@
The creation of general and multipurpose systems requires a complex and large scale data from different sources and computational resources. Geospatial data analysis can be performed using Big data tools, such as Hadoop-based frameworks, such as SpatialHadoop [^16] https://spatialhadoop.cs.umn.edu/ and HadoopGIS [^17] https://github.com/bunnyg/Hadoop-GIS, or Spark-based frameworks, including GeoTrellis [^18] http://geotrellis.github.io/ and GeoSpark [^19] https://github.com/agile-lab-dev/GeoSpark. Additionally, other proposed solutions, such as "BiGeo"[Liu2019], may also be utilized depending on the specific application requirements.
@@ 5 Applications
  • Climate Interaction & Adaptation : explores how soil data support climate modeling and mitigation.
@@ 5.1 Precision Agriculture
Ramzan et al. [Ramzan2024] proposed a two-module crop recommendation system, integrating environmental and soil parameters to guide crop selection. The first module utilizes a combination of static data, such as publicly available datasets and machine learning algorithms including K-Nearest Neighbors (KNN), Decision Tree, RF, SVM, LightGBM, CatBoost, and Adaptive Boosting (AdaBoost). This module was trained on the Kaggle “Crop Recommendation Dataset”, which contains approximately 100 samples for each of 22 crop types. The second module incorporates real-time data from Internet of Things (IoT) sensors deployed in agricultural areas of Punjab, Pakistan, focusing on six crops (wheat, rice, cotton, maize, gram, and groundnut), supplemented by manual inputs from local farmers.
@@ 5.2 Soil Fertility Analysis
Furthermore, [folorunso_exploring_2023] examined supervised machine learning algorithms, specifically, Random Forest (RF), Support Vector Machines (SVM), and Deep Neural Networks (DNN) for quantitative prediction of soil nutrient properties within precision agriculture. The research emphasized advancements in DSM methodologies and demonstrated the integration of Diffuse Reflectance Infrared Spectroscopy (DRIS) with Convolutional Neural Networks (CNN) for spectral data analysis, applying this combined approach to the LUCAS soil dataset to quantitatively estimate six soil chemical parameters.
On the other hand, [el_behairy_accurate_2024] developed a Soil Quality Index (SQI) model that uses 16 physical, chemical, and fertility soil attributes derived from laboratory analyzes, employing an Artificial Neural Network (ANN) implemented in MATLAB, without incorporating geospatial data.
In our previous study [mine_comparative_2025], we compared four machine learning algorithms (RBFN,LightGBM,DNN,XGBoost) to identify the most effective regression to predict soil quality index using the SoilGrid dataset in four European countries (Portugal, Spain, France and Italy. Based on best performance, the XGBoost was selected to generate soil quality maps.